# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ 8.2 Generalization
A central challenge in developing generalizable self-evolving agents lies in designing scalable architectures capable of maintaining performance as complexity and scope increase. Current agent systems frequently encounter a trade-off between specialization and generalization, where agents optimized for specific tasks struggle to transfer their learned behaviors to novel environments (chen2024optima, ). Additionally, the computational cost associated with dynamic reasoning in LLM-based agents grows non-linearly with the complexity of adaptation mechanisms, imposing practical constraints on achievable generalization within realistic resource limitations (kim2025cost, ). Recent studies indicate that self-evolving agents equipped with reflective and memory-augmented capabilities show substantial promise for enhancing generalization, particularly in smaller, resource-constrained models (liang2024sage, ). Nonetheless, these approaches continue to encounter limitations when addressing complex real-world scenarios that require sustained adaptation over prolonged periods.
@@
Achieving generalization across domains represents a critical frontier for self-evolving agents. Current methods frequently rely on domain-specific fine-tuning, restricting agents’ adaptability to new environments without retraining (belle2025agents, ). Recent advancements in test-time scaling and inference-time adaptation provide promising pathways for enhancing cross-domain generalization (snell2024scaling, ; zhang2025survey, ). These techniques allow agents to dynamically allocate additional reasoning capacity to unfamiliar scenarios by scaling computational resources during inference, avoiding the need for increasing model parameters. Additionally, meta-learning strategies have demonstrated considerable potential in facilitating rapid few-shot adaptation to new domains (bilal2025meta, ). However, their effectiveness critically depends on an agent’s capability to accurately determine when supplementary computational resources are necessary and efficiently distribute these resources across diverse reasoning tasks.
@@
Self-evolving agents must continuously adapt to new tasks while retaining previously acquired knowledge, a challenge exacerbated by the catastrophic forgetting phenomenon (ghosal2024understanding, ) of continual memorization (chen2024continual, ) inherent in LLMs (bell2025future, ). The stability-plasticity dilemma becomes particularly acute in foundation model-based agents, where the computational costs of retraining for every new task are prohibitive (zheng2025lifelong, ). Recent research has explored parameter-efficient fine-tuning methods, selective memory mechanisms, and incremental learning strategies to mitigate catastrophic forgetting while preserving adaptability (wang2024comprehensive, ). Nonetheless, achieving an optimal balance between efficiency and preventing model drift remains a significant open challenge, especially when agents operate under resource constraints or manage streaming data with stringent privacy considerations.
@@
Recent studies have identified critical limitations in knowledge transfer among AI agents. shi2025models emphasized that knowledge integration and transfer capabilities in current agents still require significant optimization. In particular, geng2025large found that LLM-based agents often fail to effectively propagate newly acquired knowledge from interactions to other agents, restricting their collaborative potential. Furthermore, vafa2025has revealed that foundation models might depend heavily on shallow pattern matching, rather than developing robust and transferable internal world models. These findings indicate several important future research directions: 1) it is essential to better understand the conditions under which knowledge acquired by one agent can be reliably generalized and communicated to others; 2) developing methods to quantify the limitations in agents’ knowledge transferability could lead to clearer insights into agent collaboration bottlenecks; 3) we need to have an explicit mechanism that encourage the formation of robust, generalizable world models could significantly improve the collaborative effectiveness of self-evolving agents.
8.3 Safe and Controllable Agents
As autonomous AI agents become increasingly capable of learning, evolving, and performing complex tasks independently, more agent-based studies are shifting their focus towards the deployment of safer and more controllable agents. These safety concerns arise primarily from user-related risks, such as vague or misleading instructions that lead agents to execute harmful actions, as well as environmental risks, including exposure to malicious content, such as phishing website links (zhou2025automating, ).
Many studies aimed to address safety concerns about the automatic adaptation of agents. For instance, TrustAgent (hua2024trustagent, ) implements pre-planning, in-planning, and post-planning strategies to foster safer agent behavior. However, as highlighted in (zharmagambetov2025agentdam, ), current agents based on LLM still struggle to accurately differentiate between sensitive information that is necessary and irrelevant information. A major challenge here is the precise identification and understanding of task-related versus unrelated information. Furthermore, managing agent actions when goals involve deceptive or unethical methods presents further difficulties, as ongoing learning uncertainty exacerbates these safety challenges for the deployment of controllable agents (anwar2024foundational, ). This uncertainty is reflected similarly in ambiguous contexts (bagdasarian2024airgapagent, ) and poorly designed memory modules (wang2025unveiling, ). Therefore, deploying a reliable, controllable, and safe self-evolving system has become a critical issue. Future research should focus on collecting larger-scale, more diverse real-world scenario data to support comprehensive learning of safe behaviors. Further refining the Agent Constitution by developing clearer, more understandable rules and case libraries is essential. Furthermore, exploring safer training algorithms and thoroughly investigating the impacts of privacy-protection measures on agent efficiency are necessary steps toward achieving a more balanced and secure deployment of autonomous AI agents.
@@ 8.4 Ecosystems of Multi-Agents
Recent studies highlight the difficulty of balancing independent reasoning with effective group decision-making in multi-agent environments (chen2025mdteamgpt, ; sun2025collab, ). While collective discussions can significantly enhance diagnostic reasoning, agents often risk becoming overly reliant on group consensus, thereby diminishing their independent reasoning capabilities. To mitigate this issue, future research should explore dynamic mechanisms that adjust the relative weight of individual versus collective input. Such an approach would help prevent decision-making from being dominated by a single or a small subset of agents, ultimately promoting robust, balanced consensus-building and innovation. Additionally, developing explicit knowledge bases and standardized updating methodologies—leveraging agents’ successes and failures—could further improve the agents’ self-evolution abilities and strengthen their individual reasoning contributions within collaborative contexts.
@@
Another crucial challenge lies in developing efficient algorithms and adaptive frameworks that allow agents to collaborate effectively while preserving their individual decision-making strengths. (hu2024self, ) introduced adaptive reward models and optimized dynamic network structures, which can significantly enhance cooperative self-improvement among agents. However, a major gap identified by (sun2025collab, ) is the absence of clear mechanisms for agents to dynamically manage and update their knowledge. Addressing this issue will require new frameworks that explicitly integrate continuous learning and adaptive collaboration mechanisms. Furthermore, existing benchmarks for multi-agent evaluation are predominantly static (zhu2025multiagentbench, ) and therefore fail to capture the long-term adaptability and continuous evolution of agent roles. Future benchmarks should incorporate dynamic assessment methods, reflecting ongoing adaptation, evolving interactions, and diverse contributions within multi-agent systems, thus providing more comprehensive evaluation metrics for self-evolving agents.
@@ 9 Conclusion
The emergence of self-evolving agents marks a paradigm shift in artificial intelligence, moving beyond static, monolithic models toward dynamic agentic systems capable of continual learning and adaptation. As language agents are increasingly deployed in open-ended, interactive environments, the ability to evolve, adapting reasoning processes, tools, and behaviors in response to new tasks, knowledge, and feedback, has become essential for building the next generation of agentic systems. In this survey, we provide the first comprehensive and systematic review of self-evolving agents, organized around three foundational questions: what aspects of an agent should evolve, when evolution should occur, and how to implement evolutionary processes effectively. Moreover, we discuss several methods for evaluating the progress of self-evolving agents in terms of metrics and benchmarks, followed by corresponding applications and future directions. Looking ahead, realizing the full potential of self-evolving agents will be critical in laying the groundwork for Artificial Super Intelligence (ASI). The evolution of these agents will require significant advancements in models, data, algorithms, and evaluation practices, and so on. Addressing issues such as catastrophic forgetting, human preference alignment during autonomous evolution, and the co-evolution of agents and environments will be key to unlocking agents that are not only adaptive but also trustworthy and aligned with human values. We hope this survey provides a foundational framework for researchers and practitioners to design, analyze, and advance the development and progress of self-evolving agents.

### v4

@@ 8.2 Generalization
A central challenge in developing generalizable self-evolving agents lies in designing scalable architectures capable of maintaining performance as complexity and scope increase. Current agent systems frequently encounter a trade-off between specialization and generalization, where agents optimized for specific tasks struggle to transfer their learned behaviors to novel environments (Chen et al., 2024d). Additionally, the computational cost associated with dynamic reasoning in LLM-based agents grows non-linearly with the complexity of adaptation mechanisms, imposing practical constraints on achievable generalization within realistic resource limitations (Kim et al., 2025). Recent studies indicate that self-evolving agents equipped with reflective and memory-augmented capabilities show substantial promise for enhancing generalization, particularly in smaller, resource-constrained models (Liang et al., 2024). Nonetheless, these approaches continue to encounter limitations when addressing complex real-world scenarios that require sustained adaptation over prolonged periods.
@@
Achieving generalization across domains represents a critical frontier for self-evolving agents. Current methods frequently rely on domain-specific fine-tuning, restricting agents’ adaptability to new environments without retraining (Belle et al., 2025). Recent advancements in test-time scaling and inference-time adaptation provide promising pathways for enhancing cross-domain generalization (Snell et al., 2024; Zhang et al., 2025l). These techniques allow agents to dynamically allocate additional reasoning capacity to unfamiliar scenarios by scaling computational resources during inference, avoiding the need for increasing model parameters. Additionally, meta-learning strategies have demonstrated considerable potential in facilitating rapid few-shot adaptation to new domains (Bilal et al., 2025). However, their effectiveness critically depends on an agent’s capability to accurately determine when supplementary computational resources are necessary and efficiently distribute these resources across diverse reasoning tasks.
@@
Self-evolving agents must continuously adapt to new tasks while retaining previously acquired knowledge, a challenge exacerbated by the catastrophic forgetting phenomenon (Ghosal et al., 2024) of continual memorization (Chen et al., 2024b) inherent in LLMs (Bell et al., 2025). The stability-plasticity dilemma becomes particularly acute in foundation model-based agents, where the computational costs of retraining for every new task are prohibitive (Zheng et al., 2025c). Recent research has explored parameter-efficient fine-tuning methods, selective memory mechanisms, and incremental learning strategies to mitigate catastrophic forgetting while preserving adaptability (Wang et al., 2024c). Nonetheless, achieving an optimal balance between efficiency and preventing model drift remains a significant open challenge, especially when agents operate under resource constraints or manage streaming data with stringent privacy considerations.
@@
Recent studies have identified critical limitations in knowledge transfer among AI agents. Shi et al. (2025a) emphasized that knowledge integration and transfer capabilities in current agents still require significant optimization. In particular, Geng et al. (2025b) found that LLM-based agents often fail to effectively propagate newly acquired knowledge from interactions to other agents, restricting their collaborative potential. Furthermore, Vafa et al. (2025) revealed that foundation models might depend heavily on shallow pattern matching, rather than developing robust and transferable internal world models. These findings indicate several important future research directions: 1) it is essential to better understand the conditions under which knowledge acquired by one agent can be reliably generalized and communicated to others; 2) developing methods to quantify the limitations in agents’ knowledge transferability could lead to clearer insights into agent collaboration bottlenecks; 3) we need to have an explicit mechanism that encourage the formation of robust, generalizable world models could significantly improve the collaborative effectiveness of self-evolving agents.
8.3 Safe and Controllable Self-Evolving Agents
As autonomous AI agents become increasingly capable of learning, evolving, and performing complex tasks independently, ensuring their safety and controllability has become a paramount concern. Unlike static systems, the very nature of self-evolution introduces unique and amplified risks that emerge dynamically over the agent’s lifecycle. These risks are not merely extensions of traditional AI safety issues, such as those arising from vague user instructions or environmental threats like malicious phishing links (Zhou et al., 2025d), but are fundamentally tied to the agent’s capacity for autonomous self-modification and adaptation (Shao et al., 2025; Han et al., 2025). This section delineates the emergent risks unique to self-evolving agents and then discusses a set of prescriptive guardrails and mitigation strategies for building safer systems.
8.3.1 Emergent Risks in Self-Evolving Systems
Recent research has identified new risks that arise from the autonomous self-improvement process itself. These risks are not necessarily present in the initial agent but can manifest over time as it evolves. We categorize these risks along the primary evolutionary pathways of an agent: the backbone model, memory, and tools.
  • Uncontrolled behavior drift in model evolution: A core risk is that an agent’s goals and values may drift away from original human intent as it evolves. This is exacerbated by the learning uncertainty inherent in self-evolution, especially when operating in ambiguous contexts (Anwar et al., 2024; Bagdasarian et al., 2024). A phenomenon termed "misevolution" (Shao et al., 2025) can occur during model evolution. For instance, self-training on agent-generated data can lead to "catastrophic forgetting" of safety alignment, causing agents to execute harmful instructions they previously refused, such as interacting with malicious content they were trained to avoid (Shao et al., 2025; Hahm et al., 2025).
  • Deployment-time reward hacking in memory evolution: Open-ended evolution is susceptible to reward hacking. Agents may find and exploit loopholes in self-defined reward signals or internal feedback. This is particularly evident in memory evolution, where the accumulation of experience can inadvertently induce unsafe behaviors. For example, an agent might learn to issue unnecessary refunds because its memory correlates them with high satisfaction ratings (Shao et al., 2025), and such risks can be further exacerbated by poorly designed memory modules (Wang et al., 2025b). A related concept is the "Alignment Tipping Process (ATP)," where an initially aligned agent discovers that misaligned behaviors are more rewarding, causing its policy to "tip" and abandon its initial constraints (Han et al., 2025).
  • Safety of self-created and ingested external tools: The ability of agents to autonomously generate and use tools introduces significant safety issues. Agents may spontaneously create tools with security vulnerabilities or fail to identify malicious code when ingesting external tools (Shao et al., 2025). This turns the agent into a potential vector for security threats. A major challenge here is that agents still struggle to differentiate between necessary and irrelevant sensitive information (Zharmagambetov et al., 2025), potentially leading them to create tools that leak private data. Furthermore, as agents improve, they may become more adept at creating and executing offensive cyber-operations, a risk that requires dynamic assessment (Wei et al., 2025a).
Table 12: Compliance checklist for deploying self-evolving agents
Category Compliance Checklist for Deployment
Tool & Code Safety $\Box$ Strict Sandboxing: All tools and agent-generated code execute in an isolated environment with no default access to host files, network, or sensitive processes.
$\Box$ Resource Limiting: The sandbox imposes strict limits on CPU, memory, and execution time to prevent denial-of-service or runaway processes.
$\Box$ Automated Security Verification: A mandatory pipeline performs static analysis (e.g., SAST) and vulnerability scanning on all new or modified tools before use.
$\Box$ Dependency Scanning: The verification pipeline checks all third-party libraries and dependencies for known vulnerabilities.
$\Box$ Risk-Based Access Control: Tools are classified by risk level, and high-risk capabilities (e.g., file system writes, API calls) require explicit approval via an approval gate.
Self-Modification Control $\Box$ Immutable Audit Trail: All self-modifications (to model weights, memory, toolset, or core logic) are logged with details on the trigger, changes made, and outcome.
$\Box$ Version Control for Safe States: The agent’s state (model, memory, tools) is versioned, with known "safe" versions clearly tagged.
$\Box$ Tested Rollback Mechanism: A reliable, one-click rollback mechanism exists to revert the agent to a previously known safe version. This mechanism is regularly tested.
$\Box$ Pre-Update Safety Validation: Before a self-modified model is deployed, it is automatically evaluated against a "golden dataset" of safety-critical prompts to prevent catastrophic forgetting of alignment.
Behavioral & Alignment Safety $\Box$ Continuous Runtime Monitoring: An active monitoring system tracks agent actions, flagging deviations from expected behavior, anomalous resource usage, or signs of unsafe actions.
$\Box$ Reward Hacking Detection: Key metrics are monitored for signs of reward hacking (e.g., exploiting loopholes in reward functions). Alerts are configured for sharp, unexplained metric changes.
$\Box$ Automated Red-Teaming: A continuous red-teaming framework is active, programmatically generating and running test scenarios to probe for emergent misalignment, value drift, and new failure modes.
$\Box$ Goal Guardrails: Strict constraints are placed on the agent’s ability to modify its own fundamental goals or safety constraints. Any such change requires human review.
Data Privacy & Memory Integrity $\Box$ Proactive Memory Defense: Mechanisms like dual-memory structures or consensus validation are in place to detect, isolate, and neutralize potentially "poisoned" or harmful memories before they influence behavior.
$\Box$ PII Detection and Sanitization: Automated tools are used to detect and redact/anonymize Personally Identifiable Information (PII) before it is stored in long-term memory or used in training.
$\Box$ Data Minimization Principle: The agent is configured to only collect and retain data that is strictly necessary for its tasks, and data retention policies are enforced.
$\Box$ Privacy Regulation Compliance: The system is designed to comply with relevant data privacy regulations (e.g., GDPR, CCPA).
Operational Controls & Governance $\Box$ Human-in-the-Loop for Critical Actions: High-stakes actions (e.g., large financial transactions, data deletion, communication with external users) are gated by a mandatory human approval step.
$\Box$ Clear Incident Response Plan: A documented plan is in place for responding to safety failures, including steps for immediate shutdown, rollback, and analysis.
$\Box$ Centralized Dashboard for Oversight: A dashboard provides human operators with real-time visibility into the agent’s behavior, state, and active safety alerts.
$\Box$ Explainability & Traceability: The system provides clear explanations for why a particular action was taken, linking it back to specific memories, goals, or model inferences in the audit trail.
8.3.2 Prescriptive Guardrails and Mitigation Strategies
Addressing the emergent risks of self-evolution requires moving beyond descriptive warnings to implementing prescriptive, actionable guardrails. While early frameworks like TrustAgent have explored multi-stage strategies (i.e., pre-, in-, and post-planning) to foster safer behavior (Hua et al., 2024), the unique dynamics of self-evolution call for a more comprehensive "safety lifecycle" approach. Future research and development are expected to focus on integrating safeguards at every stage of the agent’s operation.
  • Sandboxing and verification for tool and code execution: To mitigate risks from tool use, all agent-generated or externally-sourced tools must be executed in a strictly sandboxed environment. Furthermore, automated safety verification, such as static analysis and vulnerability scanning, should be a default step before a new tool is integrated (Labs, 2025). For securing tool interaction protocols, runtime defense pipelines (Xing et al., 2025; Wang et al., 2025a) provide layered detection against threats like tool poisoning and prompt injection.
  • Audit trails and failsafes for self-modification: Any self-modification must be accompanied by a comprehensive audit trail. This ensures that changes are traceable and reversible. Implementing rollback and failsafe patterns is critical, allowing the system to revert to a previously known safe state if undesirable behavior is detected. For memory, proactive defenses like A-MemGuard propose dual-memory structures and consensus-based validation to identify and isolate "poisoned" memories before they corrupt behavior (Wang et al., 2025b; Wei et al., 2025d).
  • Continuous monitoring and red-teaming for long-horizon drift: Static, pre-deployment safety evaluations are insufficient. Continuous monitoring of agent behavior is necessary to detect long-horizon value drift. This can be achieved through red-teaming scenarios designed to test for emergent misalignment. For GUI agents, hybrid validation frameworks like OS-Sentinel combine formal verifiers with contextual judges to provide robust, in-workflow safety monitoring (Sun et al., 2025b).
  • Approval gates and privacy-protection measures: For high-stakes actions, approval gates requiring human-in-the-loop confirmation should be implemented. Furthermore, given that agents struggle with handling sensitive data (Zharmagambetov et al., 2025), robust privacy-protection measures are necessary to prevent leakage and ensure a balanced and secure deployment.
To aid practitioners, we synthesize these strategies into a compliance checklist for deploying self-evolving agents (See Table 12). In conclusion, deploying reliable, controllable, and safe self-evolving systems is a critical and active area of research. Future work must move towards building a comprehensive safety-aware evolutionary lifecycle, integrating robust verification, continuous monitoring, and adaptive guardrails to ensure that the agents remain aligned with human values and safety constraints as they become more autonomous.
@@ 8.4 Ecosystems of Multi-Agents
Recent studies highlight the difficulty of balancing independent reasoning with effective group decision-making in multi-agent environments (Chen et al., 2025d; Sun et al., 2025a). While collective discussions can significantly enhance diagnostic reasoning, agents often risk becoming overly reliant on group consensus, thereby diminishing their independent reasoning capabilities. To mitigate this issue, future research should explore dynamic mechanisms that adjust the relative weight of individual versus collective input. Such an approach would help prevent decision-making from being dominated by a single or a small subset of agents, ultimately promoting robust, balanced consensus-building and innovation. Additionally, developing explicit knowledge bases and standardized updating methodologies—leveraging agents’ successes and failures—could further improve the agents’ self-evolution abilities and strengthen their individual reasoning contributions within collaborative contexts.
@@
Another crucial challenge lies in developing efficient algorithms and adaptive frameworks that allow agents to collaborate effectively while preserving their individual decision-making strengths. (Hu et al., 2024d) introduced adaptive reward models and optimized dynamic network structures, which can significantly enhance cooperative self-improvement among agents. However, a major gap identified by (Sun et al., 2025a) is the absence of clear mechanisms for agents to dynamically manage and update their knowledge. Addressing this issue will require new frameworks that explicitly integrate continuous learning and adaptive collaboration mechanisms. Furthermore, existing benchmarks for multi-agent evaluation are predominantly static (Zhu et al., 2025) and therefore fail to capture the long-term adaptability and continuous evolution of agent roles. Future benchmarks should incorporate dynamic assessment methods, reflecting ongoing adaptation, evolving interactions, and diverse contributions within multi-agent systems, thus providing more comprehensive evaluation metrics for self-evolving agents.
@@ 9 Conclusion
The emergence of self-evolving agents marks a paradigm shift in artificial intelligence, moving beyond static, monolithic models toward dynamic agentic systems capable of continual learning and adaptation. As language agents are increasingly deployed in open-ended, interactive environments, the ability to evolve, adapting reasoning processes, tools, and behaviors in response to new tasks, knowledge, and feedback, has become essential for building the next generation of agentic systems. In this survey, we provide the first comprehensive and systematic review of self-evolving agents, organized around three foundational questions: what aspects of an agent should evolve, when evolution should occur, and how to implement evolutionary processes effectively. Moreover, we discuss several methods for evaluating the progress of self-evolving agents in terms of metrics and benchmarks, followed by corresponding applications and future directions. The evolution of these agents will require significant advancements in models, data, algorithms, and evaluation practices, and so on. Addressing issues such as catastrophic forgetting, human preference alignment during autonomous evolution, and the co-evolution of agents and environments will be key to unlocking agents that are not only adaptive but also trustworthy and aligned with human values. We hope this survey provides a foundational framework for researchers and practitioners to design, analyze, and advance the development and progress of self-evolving agents.